### v1

@@ 1 Introduction
Directed searches lie in-between these two extremes, with known sky positions, but unknown spin frequency. This type of searches include interesting spots such as the galactic centre or the globular cluster Terzan 5 (Piccinni et al., 2020; Dergachev et al., 2019; Abbott et al., 2022d), thought to host a large number of neutron stars, young supernova remnants (Ming et al., 2019; Papa et al., 2020; Abbott et al., 2019c, d; Millhouse et al., 2020; Lindblom & Owen, 2020; Abbott et al., 2021e, 2022d, 2022a; Owen et al., 2022; Liu & Zou, 2022; Ming et al., 2022), and low-mass X-ray binaries such as Scorpius X-1 (Zhang et al., 2021; Whelan et al., 2023).
@@
Guided by this study, in the quest to detect a signal, a number of searches have been performed targeting these objects on more and more sensitive data. We continue in this vein by extending to higher frequencies and to higher sensitivity previous searches for emission from G347.3 (Ming et al., 2019; Papa et al., 2020; Ming et al., 2022). We also search for emission from Vela Jr, with a different method, a broader frequency-evolution model, and a different data set than the latest search for emission from this object, reported by Abbott et al. (2022a). In spite of us using data about a factor 1.6 less sensitive, our sensitivity to Vela Jr. is less than 20% worse than that of Abbott et al. (2022a).
@@ 4.1 Semi-coherent searches
The $\mathcal{F}_{i}$-statistic is a coherent combination of the data from all detectors in the $i$-th segment, whereas the summation 5 of the detection statistic values is not. This is the reason why this type of search is called “semi-coherent”.
@@
The first stage is the most computationally expensive because the entire parameter space is searched. Based on the detection statistic values, the most promising candidates are identified and only those are passed on to the second stage. In the second stage an appropriate volume of parameter space is searched around the nominal candidate parameters, and if the outcome is consistent with what expected from a signal, the candidate passes on to the third stage, else it is discarded. In total we have four stages, from 0 to 3.
@@ 5 Hierarchical follow up searches
$\textrm{if}~{}R^{\textrm{(a)}}<R^{\textrm{(a-thr)}}\longrightarrow{\textrm{candidate discarded}},$ (8)
where
$R^{\textrm{(a)}}:={{2\overline{\mathcal{F}}^{\textrm{(a)}}-4}\over{{2\overline{\mathcal{F}}^{\textrm{(0)}}-4}}},$ (9)
and the superscript “0” indicates that the detection statistic value comes from the original $\mathrm{Einstein@Home}$ search. For signals the expected value of $R^{\textrm{(a)}}\propto T_{\textrm{\mbox{\tiny{coh}}}}^{(a)}/T_{\textrm{\mbox{\tiny{coh}}}}^{(0)}$, so we expect it to increase as a signal candidate is investigated with longer and longer $T_{\textrm{\mbox{\tiny{coh}}}}$.
The threshold values are determined based on the distributions of $R^{\textrm{(a)}}$ for our target signals, which are shown in Figs. 2, 3 and 4. In particular, we fit the normalized histograms ($n_{i}/N_{tot}$) with a Gaussian and set the threshold such that the false dismissal probability is always $\lambda_{\text{FD}}\ll 0.1\%$, as shown in Table 5. We experiment in constructing different discriminators, also using the transient- and line-robust statistic ${\hat{\beta}}_{{{\mathrm{{S}}}/{\mathrm{{G}}{\mathrm{{L}}}{\mathrm{t{L}}}}}}$ but find that Eq. 9 defines the most efficient detection statistic to identify the candidates for the next stage. With “most efficient” here we mean that at fixed false dismissal it yields the lowest false alarm.
@@
5.1 Banding and identify disturbed ones
@@ 5.2 Clustering
The mean mismatch of our searches is about 13% which means that neighboring templates are not entirely independent. No matter whether it’s a disturbance or a signal, both can generate multiple adjacent detection-statistic results that are prominent enough to make our top list. However, we don’t need to follow up all of them, because their origin is the same. To avoid wasting computing resources on these “repetitive” results, we use a clustering procedure that identifies results that originate from the same underlying cause and that allows us to consider them as a single candidate for the purpose of follow-ups. Different clustering have been developed over time to process the large Einstein@Home top lists (A. Singh & Walsh, 2017; Beheshtipour & Papa, 2020, 2021). Here we consider our most recent and effective one (Steltner et al., 2022a). We refer to these output of the clustering procedure as “candidates”.
@@
$R^{\textrm{(a)}}$ $\lambda_{\text{FD}}$ $N_{\textrm{in}}$ $N_{\textrm{out}}$ $R^{\textrm{(a)}}$ $\lambda_{\text{FD}}$ $N_{\textrm{in}}$ $N_{\textrm{out}}$ $R^{\textrm{(a)}}$ $\lambda_{\text{FD}}$ $N_{\textrm{in}}$ $N_{\textrm{out}}$
@@ 5.5 Stage 3
Only two candidates survive the Stage 3 cut, from the Vela Jr. search, as shown in Fig. 2. One is at a frequency $\approx$ 42 Hz with a $37{2\mathcal{F}_{r}}=$ 37 and the other is at a frequency $\approx$ 100 Hz with a $30{2\mathcal{F}_{r}}=$ 30. However the corresponding ${\hat{\beta}}_{{{\mathrm{{S}}}/{\mathrm{{G}}{\mathrm{{L}}}{\mathrm{t{L}}}}r}}$ values are $-18$ and $-23$ respectively, making them the two lowest values of all 180, 000 Vela Jr. candidates followed-up in Stage 3, suggesting that a disturbance in one of the two detectors is the likely cause of the elevated values of $2\mathcal{F}_{r}$. This is confirmed by inspecting the average power spectral density from the O3a data for the detectors. We find large spectral features right around the candidates’ frequencies in the H1 detector.
@@ 7 Discussion and Conclusions
Targeted searches for emission from known pulsars have excluded much smaller ellipticities (Abbott et al., 2019a, 2021a; Ashok et al., 2021; Abbott et al., 2022b), so one could argue that we already know that neutron stars are extremely spherical. Pulsars are however generally much older than the young neutron stars targeted here. Vela Jr. and G347.3 are much closer to their birth than typical pulsars, and that birth process is a catastrophic event unlikely to be spherically symmetric. It is not unlikely that the newborn neutron star presents some leftover non-axisymmetry (Janka et al., 2008). This is however expected to reduce in time as the star cools down (Gnedin et al., 2001; Haensel et al., 1990), producing a population of very spherical older pulsars.

### v2

@@ 1 Introduction
Directed searches lie in-between these two extremes, with known sky positions, but unknown spin frequency. This type of searches include interesting spots such as the galactic centre or the globular cluster Terzan 5 (Piccinni et al., 2020; Dergachev et al., 2019; Abbott et al., 2022d), thought to host a large number of neutron stars, young supernova remnants (Ming et al., 2019; Papa et al., 2020; Abbott et al., 2019c, d; Millhouse et al., 2020; Lindblom & Owen, 2020; Abbott et al., 2021e, 2022d, 2022a; Owen et al., 2022; Liu & Zou, 2022; Ming et al., 2022; Wang & Riles, 2024), and low-mass X-ray binaries such as Scorpius X-1 (Zhang et al., 2021; Whelan et al., 2023).
@@
Guided by this study, in the quest to detect a signal, a number of searches have been performed targeting these objects on more and more sensitive data. We continue in this vein by extending to higher frequencies and to higher sensitivity previous searches for emission from G347.3 (Ming et al., 2019; Papa et al., 2020; Ming et al., 2022). We also search for emission from Vela Jr, with a different method, a broader frequency-evolution model, and a different data set than the latest search for emission from this object, reported by Abbott et al. (2022a). Despite of us using data about a factor 1.6 less sensitive, our sensitivity to Vela Jr. is less than 20% worse than that of Abbott et al. (2022a).
@@ 4.1 Semi-coherent searches
The $\mathcal{F}_{i}$-statistic is a coherent combination of the data from all detectors in the $i$-th segment, whereas the summation of the detection statistic values in Eq. 5 is not. This is the reason why this type of search is called “semi-coherent”.
@@
The first stage is the most computationally expensive because the entire parameter space is searched. Based on the detection statistic values, the most promising candidates are identified and only those are passed on to the second stage. In the second stage an appropriate volume of parameter space is searched around the nominal candidate parameters, and if the outcome is consistent with what expected from a signal, the candidate passes on to the third stage, else it is discarded. In total we have four stages, numbered from 0 to 3.
@@ 5 Hierarchical follow up searches
$\textrm{if}~{}R^{\textrm{(a)}}<R^{\textrm{(a)}}_{\textrm{thr}}\longrightarrow{\textrm{candidate discarded}},$ (8)
where $a$ is the stage index, and
$R^{\textrm{(a)}}:={{2\overline{\mathcal{F}}^{\textrm{(a)}}-4}\over{{2\overline{\mathcal{F}}^{\textrm{(0)}}-4}}}.$ (9)
The superscript “0” indicates that the detection statistic value comes from the original $\mathrm{Einstein@Home}$ search. For signals the expected value of $R^{\textrm{(a)}}\propto T_{\textrm{\mbox{\tiny{coh}}}}^{(a)}/T_{\textrm{\mbox{\tiny{coh}}}}^{(0)}$ (see Eq. 6), so we expect it to increase as a signal candidate is investigated with longer and longer $T_{\textrm{\mbox{\tiny{coh}}}}$.
The threshold values $R^{\textrm{(a)}}_{\textrm{thr}}$ are determined based on the distributions of $R^{\textrm{(a)}}$ for our target signals, which are shown in Figs. 2, 3 and 4. In particular, we fit the normalized histograms with a Gaussian and set the threshold such that the false dismissal probability is always $\lambda_{\text{FD}}\ll 0.1\%$, as shown in Table 5. We experiment in constructing different discriminators, also using the transient- and line-robust statistic ${\hat{\beta}}_{{{\mathrm{{S}}}/{\mathrm{{G}}{\mathrm{{L}}}{\mathrm{t{L}}}}}}$ but find that Eq. 9 defines the most efficient detection statistic to identify the candidates for the next stage. With “most efficient” here we mean that at fixed false dismissal it yields the lowest false alarm.
@@
5.1 Banding and identification of disturbed bands
@@ 5.2 Clustering
The mean mismatch of our searches is about 13% which means that neighboring templates are not entirely independent. No matter whether it is a disturbance or a signal, both can generate multiple adjacent detection-statistic results that are prominent enough to make them into our top list. However, we don’t need to follow up all of them, because their origin is the same. To avoid wasting computing resources on these “repetitive” results, we use a clustering procedure that identifies results that originate from the same underlying cause and that allows us to consider them as a single candidate for the purpose of follow-ups. Different clustering have been developed over time to process the large Einstein@Home top lists (A. Singh & Walsh, 2017; Beheshtipour & Papa, 2020, 2021). Here we consider our most recent and effective one (Steltner et al., 2022a). We refer to these output of the clustering procedure as “candidates”.
@@
$R^{\textrm{(a)}}_{\textrm{thr}}$ $\lambda_{\text{FD}}$ $N_{\textrm{in}}$ $N_{\textrm{out}}$ $R^{\textrm{(a)}}_{\textrm{thr}}$ $\lambda_{\text{FD}}$ $N_{\textrm{in}}$ $N_{\textrm{out}}$ $R^{\textrm{(a)}}_{\textrm{thr}}$ $\lambda_{\text{FD}}$ $N_{\textrm{in}}$ $N_{\textrm{out}}$
@@ 5.5 Stage 3
Only two candidates survive the Stage 3 cut, from the Vela Jr. search, as seen in Fig. 2. One is at a frequency $\approx$ 42 Hz with a $37{2\mathcal{F}_{r}}=$ 37 and the other is at a frequency $\approx$ 100 Hz with a $30{2\mathcal{F}_{r}}=$ 30. However the corresponding ${\hat{\beta}}_{{{\mathrm{{S}}}/{\mathrm{{G}}{\mathrm{{L}}}{\mathrm{t{L}}}}r}}$ values are $-18$ and $-23$ respectively, making them the two lowest values of all 180, 000 Vela Jr. candidates followed-up in Stage 3, suggesting that a disturbance in one of the two detectors is the likely cause of the elevated values of $2\mathcal{F}_{r}$. This is confirmed by inspecting the average power spectral density from the O3a data for the detectors. We find large spectral features right around the candidates’ frequencies in the H1 detector.
@@ 7 Discussion and Conclusions
Targeted searches for emission from known pulsars have excluded much smaller ellipticities (Abbott et al., 2019a, 2021a; Ashok et al., 2021; Abbott et al., 2022b), so one could argue that we already know that neutron stars are extremely axis-symmetric. Pulsars are however generally much older than the young neutron stars targeted here. Vela Jr. and G347.3 are much closer to their birth than typical pulsars, and that birth process is a catastrophic event unlikely to be spherically symmetric. It is not unlikely that the newborn neutron star presents some leftover non-axisymmetry (Janka et al., 2008). This is however expected to reduce in time as the star cools down (Gnedin et al., 2001; Haensel et al., 1990), producing a population of very spherical older pulsars.